# Uncertainty Quantification for Prior-Data Fitted Networks using Martingale Posteriors

### v1

@@ B.2 Data generation for diffusion process
    Output transformation: The input is centered to be within $[-2.5,2.5]$, and the output values are normalized based on their min-max range for potential use in quantile-based learning.
B.3 Data generation for quantile regression
For the data in Section 4.4, we sample values $x\sim U(0,1)$ and generate the corresponding outcome via $y\sim\mathcal{N}(\sin(3x),x^{2})$. A visualization of the data set is given in Figure 6.
Figure 6: Depiction of the data used in Section 4.4.
B.4 Prompt for LLMs
@@
B.5 Computational environment
All computations were performed on a user PC with Intel(R) Core(TM) i7-8665U CPU @ 1.90GHz, 8 cores, 16 GB RAM using Python 3.8, R 4.2.1, and TensorFlow 2.10.0. Run times of each experiment do not exceed 24 hours.
@@ Appendix C Further discussion
Another open problem is joint posterior inference for a collection of parameters $\theta=\{\theta(x)\colon x\in\mathcal{X}\}$. Although our proposed inference procedure can be repeated for many values of $x$, the resulting posteriors are disconnected. For obtaining a full joint posterior $\Pi(\theta|\mathcal{D}_{n})$, the distribution of the features $x_{1:n}$ can no longer be ignored. [8] proposed a general joint update of the PPDs for all values of $x$ simultaneously. However, this general update is intractable, and the heuristic simplifications proposed by [8] are neither particularly simple nor theoretically justified. There are several potential ways forward. A simple practical solution would be to specify a joint distribution that combines the individual posteriors $P_{\infty,x_{1}},\dots,P_{\infty,x_{K}}$ in a plausible way; for example, using a multivariate Gaussian copula with covariance kernel depending on the distance between values $x_{i}\neq x_{j}$. We expect such a heuristic correction to work reasonably well in many applications. A more sophisticated alternative was recently proposed by [13] and involves nonparametric estimation of the implicit dependence between PPDs by a nonparametric vine copula.

### v2

@@ B.2.6 Data Generation for Diffusion Process
    Output transformation: The input is centered to be within [2.5,2.5][-2.5,2.5], and the output values are normalized based on their min-max range for potential use in quantile-based learning.
B.2.7 Data Generation for Quantile Regression
For the data in Section B.2.2, we sample values $x\sim U(0,1)$ and generate the corresponding outcome via $y\sim\mathcal{N}(\sin(3x),x^{2})$. A visualization of the dataset is given in Figure 10.
Figure 10: Depiction of the data used in Section B.2.2.
B.2.8 Prompt and Details for LLM Experiment
@@
B.3 Computational Environment
All computations were performed on a user PC with Intel(R) Core(TM) i7-8665U CPU @ 1.90GHz, 8 cores, 16 GB RAM, using Python 3.8, R 4.2.1, and TensorFlow 2.10.0. Run times of each experiment do not exceed 24 hours.
@@ Appendix C Further Discussion
Another open problem is joint posterior inference for a collection of parameters $\theta=\{\theta(x)\colon x\in\mathcal{X}\}$. Although our proposed inference procedure can be repeated for many values of $x$, the resulting posteriors are disconnected. For obtaining a full joint posterior $\Pi(\theta|\mathcal{D}_{n})$, the distribution of the features $x_{1:n}$ can no longer be ignored. Fong et al., (2023) proposed a general joint update of the PPDs for all values of $x$ simultaneously. However, this general update is intractable, and the heuristic simplifications proposed by Fong et al., (2023) are neither particularly simple nor theoretically justified. There are several potential ways forward. A simple practical solution would be to specify a joint distribution that combines the individual posteriors $P_{\infty,x_{1}},\dots,P_{\infty,x_{K}}$ in a plausible way; for example, using a multivariate Gaussian copula with covariance kernel depending on the distance between values $x_{i}\neq x_{j}$. We expect such a heuristic correction to work reasonably well in many applications. A more sophisticated alternative was recently proposed by Huk et al., (2024) and involves nonparametric estimation of the implicit dependence between PPDs by a nonparametric vine copula.